A particle moves along x-axis under the action of a position dependent force f =(5x^2- 2X)N . work done by force on the particle when it moves from origin to X =3 is ?

Grade:11

1 Answers

Susmita
5 years ago
Work done
dW=fdx=(5x2-2x)dx
For W ,integrate both side.The limit of x is 0 to 3.
W= \int_{0}^{3} (5x^2 -2x)dx
    W=5[\frac{x^3}{3}]{_{0}^{3}}-2[\frac{x^2}{2}]{_{0}^{3}}
     W=5[\frac{27}{3}]-2[\frac{9}{2}]
W=45-9=36Nm
Hope this helps